Suppose you take out a 30-year mortgage of $100,000 with a fixed interest rate of 5 percent. You must make 360 equal monthly payments. Write an equation that defines what the payment is?


1
Expert's answer
2021-06-21T12:19:30-0400

given,

p = $100,000

r = 5%

m = 12

n = 30 

monthly payment equation:

"\frac{P\times\frac{r}{m}}{1-(1+\frac{r}{m})^{-n\times m}}"


"\frac{10000\times\frac{0.05}{12}}{1-(1+\frac{0.05}{12})^{-12\times 30}}"